OpenStudy (anonymous):

Solve 19.9 = 9y - 8.

OpenStudy (calculusxy):

\(19.9 = 9y - 8\) \(19.9 + 8 = 9y + 8 - 8\) \(27.9 = 9y\) \(\large \frac{27.9}{9} = \frac{9y}{9}\) \(3.1 = y\)